The proportion of defective items in a certain manufacturing process is 0.20. If 10
items are selected from the process, Find
a. The expected number of good items
b. Probability that none is defective
c. Probability that at least 3 are defective
1
Expert's answer
2020-04-02T13:12:12-0400

Let "X=" the number of defective items: "X\sim Bin(n, p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "p=0.20, n=10"


a. The expected number of good items


"E(X^C)=n(1-p)=10(1-0.2)=8"

b. Probability that none is defective 


"P(X=0)=\binom{10}{0}0.2^0(1-0.2)^{10-0}=0.8^{10}=0.1073741824"

c. Probability that at least 3 are defective


"P(X\geq3)=1-P(X=0)-P(X=1)-P(X=2)="

"=1-\binom{10}{0}0.2^0(1-0.2)^{10-2}-\binom{10}{1}0.2^1(1-0.2)^{10-1}-"

"-\binom{10}{2}0.2^2(1-0.2)^{10-2}=1-0.1073741824-"

"-0.268435456-0.301989888=0.3222004736\approx"

"\approx0.3222"